# Tractable Equilibrium Computation in Markov Games through Risk Aversion

### v1

@@ Appendix D Proof of Theorem 5
The remainder of the proof will focus on controlling (33). Applying [5, Lemma 17] over all $(h,s,\bm{a})\in[H]\times{\mathcal{S}}\times\mathcal{A}$, we achieve the union bound that with probability at least $1-\delta$,
@@
Appendix E Experiment Details
@@ D $\begin{matrix}370\\ 200\end{matrix}$ $\begin{matrix}10\\ 370\end{matrix}$
We use the penalty functions KL and reverse KL with the regularizers log barrier and negative entropy, as mentioned in Table 1, for our experiments. To be precise, we consider the following two experimental setups: (a) For $j=1,2$, we let $\nu_{j}(p)=-\sum_{i}\log(p_{i})$ and $D_{j}(p,q)=(1/\tau_{j})\mathrm{KL}(p,q)$. (b) For $j=1,2$, we let $\nu_{j}(p)=\sum_{i}p_{i}\log(p_{i})$ and $D_{j}(p,q)=(1/\tau_{j})\mathrm{KL}(q,p)$. We note that the parameters $\tau_{j}^{-1}$ play the same role as $\xi_{j}$’s (formalized in Corollaries 5.1 and 5.2), i.e., the players take more risk neutral decisions as $\tau_{j}\to 0$. With the perfect information, we use the vanilla projected gradient descent with constant stepsizes [6] as the no-regret algorithm to arrive at the CCE $\sigma$ of the four player game. We use consistent stepsizes between $10^{-4}$ and $10^{-3}$ for all our runs iterating through $10^{4}$ steps of gradient descent. We notice $2\%$ deviations in sup-norm for the algorithm policies in about 20 runs.
E.2 Markov Games

### v2

@@ Appendix C Proof of Theorem 5
The remainder of the proof will focus on controlling (39). Applying [3, Lemma 17] over all $(h,s,\bm{a})\in[H]\times{\mathcal{S}}\times\mathcal{A}$, we achieve the union bound that with probability at least $1-\delta$,
@@
Appendix D Experiment Details
@@ D $\begin{matrix}370\\ 200\end{matrix}$ $\begin{matrix}10\\ 370\end{matrix}$
We use the penalty functions KL and reverse KL with the regularizers log barrier and negative entropy, as mentioned in Table 1, for our experiments. To be precise, we consider the following two experimental setups: (a) For $j=1,2$, we let $\nu_{j}(p)=-\sum_{i}\log(p_{i})$ and $D_{j}(p,q)=(1/\tau_{j})\mathrm{KL}(p,q)$. (b) For $j=1,2$, we let $\nu_{j}(p)=\sum_{i}p_{i}\log(p_{i})$ and $D_{j}(p,q)=(1/\tau_{j})\mathrm{KL}(q,p)$. We note that the parameters $\tau_{j}^{-1}$ play the same role as $\xi_{j}$’s (formalized in Corollaries 6.1 and 6.2), i.e., the players take more risk neutral decisions as $\tau_{j}\to 0$. With the perfect information, we use the vanilla projected gradient descent with constant stepsizes [6] as the no-regret algorithm to arrive at the CCE $\sigma$ of the four player game. We use consistent stepsizes between $10^{-4}$ and $10^{-3}$ for all our runs iterating through $10^{4}$ steps of gradient descent. We notice $2\%$ deviations in sup-norm for the algorithm policies in about 20 runs.
D.2 Markov Games